. given the function h described by h(x) = x + 13, find each of the following.
h(0) = _____
h(-18) = _____
h(-7) = ____
h(3) = _____
h(c + 6) = _____

Answer by jim_thompson5910(28476)

h(0) = 0+13 = 13

h(-18) = -18 + 13 = -5

h(-7) = -7 + 13 = 6

h(3) = 3 + 13 = 16

h(c + 6) = (c+6) + 13 = c + 19


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim